How much work in Joules is required to expand a gas from a volume of
234 mL to a volume of 878 mL by exerting a constant pressure of 1.27 atm?
1
Expert's answer
2019-11-27T07:12:42-0500

Work,"W=P*dV"

Where pressure P is (1.27*101325) pascals and dV is change in volume.

dV = (878ml-234ml) *10^-5= 6.44*10^-3.

W=128682.75pa*6.44*10^-3.

W=828.72joules